# Chain-of-Thought Reasoning Improves Context-Aware Translation with Large Language Models

### v1

@@ Abstract
This paper assesses the capacity of large language models (LLMs) to translate texts that include inter-sentential dependencies. We use the English-French DiscEvalMT benchmark (Bawden et al., 2018) with pairs of sentences containing translation challenges either for pronominal anaphora or for lexical cohesion. We evaluate 12 LLMs from the DeepSeek-R1, GPT, Llama, Mistral and Phi families on two tasks: (1) distinguishing a correct translation from a wrong but plausible one; (2) generating a correct translation. We compare prompts that encourage chain-of-thought reasoning with those that do not. The best models take advantage of reasoning and reach about 90% accuracy on the first task, and COMET scores of about 92% on the second task, with GPT-4, GPT-4o and Phi standing out. Moreover, we observe a “wise get wiser” effect: the improvements through reasoning are positively correlated with the scores of the models without reasoning.
Keywords: Context-aware MT, Translation with LLMs, Chain-of-thought reasoning in LLMs.
@@
Shabnam Ataee1, Andrei Popescu-Belis1,2
@@
{shabnam.ataee, andrei.popescu-belis}@heig-vd.ch
@@ 1.  Introduction
Large language models (LLMs) have shown impressive capacities for translation, among many other tasks, but their translations are still not perfect (Kocmi et al., 2024; Cui et al., 2025). To increase quality, we examine whether chain-of-thought (CoT) reasoning may help to improve referential and lexical cohesion across sentences. Indeed, when prompted to generate a step-by-step explanation of their answers, LLMs improve substantially on many tasks (Wei et al., 2022).
We hypothesize that CoT can improve the translation of elements that maintain coherence across sentences by making translation decisions explicit. To test our hypothesis, we use the DiscEvalMT English-French test suites for pronominal anaphora and lexical consistency designed by Bawden et al. (2018), in two settings. First, we ask the LLM to select the correct translation among the two contrastive alternatives present in the data. Second, we ask the LLM to translate each test sentence and score the result against its reference translation. In both settings, we demonstrate the advantages of CoT prompting, though only for the largest LLMs, which improve their performance when encouraged to reason before translating.
The paper is organized as follows. We review related work in Section 2, and present the benchmark data and the evaluation metrics in Section 3. The tested LLMs are listed in Section 4: they include 4 GPT proprietary models from OpenAI and 8 open-weight models. The prompts and the results for the contrastive evaluation task (i.e., selecting the correct translation) are presented in Section 5, while those for the translation task are in Section 6. The main contributions are:
@@
  • We show that the translation task is reliably scored, as BLEU, chrF, BERTscore, and COMET scores vary in the same directions.
  • We demonstrate that CoT prompting improves coherence, but only for the best models. The improvement is positively correlated with the initial score of each LLM, showing that powerful models have a better capacity to leverage reasoning, i.e. a “wise get wiser” effect.
@@ 2.  Related Work
The translation capabilities of multilingual large language models (LLMs) have attracted growing attention since 2023, and conversational LLMs are now often used for translation tasks – around 4.5% of non-professional interactions with ChatGPT (Chatterji et al., 2025). Therefore, studies evaluating the translation capability of LLMs are of high importance, all the more that this is not tested by common benchmarks for LLMs. LLMs have been included since 2023 in the evaluations carried out at the Workshop on Machine Translation (Kocmi et al., 2023, 2024), and in 2025, the majority of submissions to the general task were LLM-based (Kocmi et al., 2025). At the same time, the first studies of LLMs for MT also include quantitative evaluations (Vilar et al., 2023; Zhang et al., 2023; Hendy et al., 2023). Other studies are focused on evaluation only: for instance, Bawden and Yvon (2023) demonstrate the strong MT performance of the BLOOM LLM, particularly in few-shot settings and for high-resource language pairs. Fine-tuning of LLMs for document-level MT was found to be moderately beneficial (Wu et al., 2024).
@@
Word-sense disambiguation (WSD) for translation is one of the challenges on which LLMs have been evaluated. Test suites related to WSD include DiscEvalMT (Bawden et al., 2018) which we use in this paper and describe below, as well as ContraWSD (Rios Gonzales et al., 2017), with about 7,000 sentences with reference translations (for De-En and De-Fr) and a smaller test suite to score word translations (Rios et al., 2018). The DiBiMT benchmark with about 600 examples, for translation from English to five other languages, was proposed by Campolungo et al. (2022) but remains private and systems must be submitted to its owners. DiBiMT was used by Iyer et al. (2023) to score the capacity of LLMs for WSD, either out-of-the-box, or after fine-tuning, or with in-context learning.
@@
Similarly, the capacity to translate potentially ambiguous pronouns can be tested using test suites such as DiscEvalMT (Bawden et al., 2018), ContraPRO (Müller et al., 2018; Lopes et al., 2020), or PROTEST (Guillou and Hardmeier, 2016), which was used at WMT 2018 (Guillou et al., 2018). A test suite for En-Ru verb generation in case of ellipses was designed by Voita et al. (2019). A review of scores reached on the DiscEvalMT data by various context-aware MT systems appears below in Section 5.3.
@@
The correct translation of pronouns or ambiguous words is often related to the more general capacity of leveraging inter-sentential dependencies and generating a contextually-correct translation, reviewed in several studies (Popescu-Belis, 2019; Maruf et al., 2021; Jin et al., 2023; Castilho and Knowles, 2025). Document-level translation and its evaluation are important because MT outputs may appear competitive with human translations at the sentence level, but not at the document level (Läubli et al., 2018). An evaluation of GPT-3.5 and GPT-4 on document-level translation (Wang et al., 2023) found that they surpassed non-LLM systems in human ratings. Karpinska and Iyyer (2023) found that LLMs translate better entire paragraphs than individual sentences, in the case of literary translation. Beyond direct translation, interaction was a key element in the WSD process proposed by Pilault et al. (2023), using questions and answers as a chain-of-thought input to an LLM.
@@
Recent work has explored how reasoning can improve LLM-based translation, based on the seminal work by Wei et al. (2022), who introduced chain-of-thought (CoT) prompting and showed that guiding models through intermediate steps improves performance on complex tasks. Liu et al. (2025) argue that reasoning enhances translation by improving coherence, cultural alignment, and self-reflection. He et al. (2025) present R1-T1, which uses expert CoT templates and reinforcement learning to encourage inference-time reasoning for translation. Ye et al. (2025) find that reasoning models outperform standard LLMs on semantically complex and domain-specific MT, especially for longer texts.
@@
Traditional metrics such as BLEU and chrF (Papineni et al., 2002; Popović, 2015), based on surface-level overlap, offer only limited insight into discourse-level quality. Embedding-based metrics such as BERTScore (Zhang et al., 2020), and trained metrics such as COMET (Rei et al., 2020), improve upon this by modeling semantic similarity and are frequently used, as we do here. LLMs have been put to use for MT evaluation: Kocmi and Federmann (2023) showed that GPT-based metrics reach state-of-the-art correlations with human judgments on WMT benchmarks. The benefits of LLM reasoning for this task are still under investigation: Larionov et al. (2025) found that reasoning-enabled LLMs such as DeepSeek-R1 and OpenAI’s o3-mini do not always outperform their non-reasoning counterparts in terms of alignment with human judgments of quality.
@@ 3.1.  Benchmark Data
This study uses the DiscEvalMT benchmark (Bawden, 2018; Bawden et al., 2018), designed to evaluate challenges in English-to-French MT that arise from two types of inter-sentential dependencies: pronominal anaphora and lexical cohesion. [^1] https://github.com/rbawden The test cases are manually constructed, and aim to test whether MT systems can make contextually-appropriate translation decisions.
EN context: The buildings will be finished next week.
EN current: Soon they will be full of new residents.
FR translation of context (given): Les bâtiments seront terminés la semaine prochaine.
FR translation #1 of current (correct): Ils seront bientôt pleins de nouveaux résidents.
FR translation #2 of current (incorrect): Elles seront bientôt pleines de nouveaux résidents.
Table 1: A contrastive test item for pronominal anaphora from DiscEvalMT (Bawden et al., 2018).
Each test item consists of two English sentences (‘context’ and ‘current’), a French translation of the first, and two alternative French translations of the second: one is contextually appropriate, preserving coherence across the two sentences, while the other introduces a discourse-level error; together, they form a contrastive pair among which a system should distinguish the correct translation.
An item for testing the translation of pronominal anaphora is shown in Table 1. The first candidate has correct gender agreement between the pronoun ‘ils’ (masculine plural) and the antecedent ‘les bâtiments’ (the translation of ‘buildings’), while the second one has wrong agreement (pronoun ‘elles’, feminine plural).
To control for chance, test items come in couples, which have a slight difference in the translation of context sentences, but have the two same alternative translations of the second source sentence. However, in one item the first alternative is correct, while in the second item, it is the second one. For instance, the item shown in Table 1 is accompanied in the dataset by another one, where the translation of ‘buildings’ is by the feminine noun ‘résidences’, which now makes the second translation correct.
For lexical choice, each item contains a word in the context sentence which can be translated in several ways, either because it is polysemous, or because the translations are synonymous. The correct translation of the current sentence is the one which translates this word with the one used in the reference translation of the first sentence.
The anaphora test set includes 200 items (100 couples), each with two candidates. Similarly, the lexical choice test set includes 200 items with two candidates each. We selected the first half of the data for development of the prompts, and keep the second half for final testing.
@@ 3.2.  Metrics
In the first part of our study (Section 5), the task of the LLM is to distinguish the correct from the wrong translation (contrastive task), while in the second part (Section 6) the task of the LLM is to translate the English current sentence into French.
@@
We score the first task using the number of times the correct translation was identified. The LLMs are prompted to execute the task by outputting their response either as ‘Choice: (1)’ or as ‘Choice: (2)’. This can be preceded by reasoning in case of CoT prompting. Our evaluation script is tolerant to minor formatting differences in the LLM’s output, and responses are accepted as correct as long as they clearly indicate the right option (1 or 2); if they cannot be parsed, they are counted as incorrect. The mean accuracy score (ACC) is the average over all the answers.
@@
To avoid systematic bias in favor of either the first or the second (last) sentence, we present each item twice, once with the correct option being first, and once second. Accordingly, we also measure consistency, i.e. the sensitivity of the model to the position of the correct translation among the two options. Ideally, the model’s response should not depend on the position, therefore we define inconsistency (INC) as the normalized difference between the accuracy score when the correct option is presented first ($\text{ACC}_{\text{correct=1}}$) and the accuracy scores when it is presented second ($\text{ACC}_{\text{correct=2}}$). A lower inconsistency indicates a more reliable LLM.
$\text{INC}=\frac{|\text{ACC}_{\text{correct=1}}-\text{ACC}_{\text{correct=2}}|}{\text{ACC}_{\text{correct=1}}+\text{ACC}_{\text{correct=2}}}$
@@
For the translation task, we provide the LLM with the two EN source sentences, as well as the FR translation of the context sentence, and ask it to translate the second EN sentence (again, possibly outputting reasoning before it). We parse the output to isolate the translation, and score it by comparing it to the correct translation using four metrics: BLEU, chrF, BERTscore and COMET. The first two capture surface-level overlap, while BERTScore and COMET assess semantic adequacy and fluency using embeddings. All metrics are computed using their official Python implementations available via the sacrebleu [^2] https://github.com/mjpost/sacrebleu, bert-score [^3] https://github.com/Tiiiger/bert_score, and unbabel-comet [^4] https://github.com/Unbabel/COMET packages. BLEU and chrF are calculated with the default configurations of sacrebleu. BERTScore is computed using the default multilingual model bert-base-multilingual-cased, with the language parameter set to French (lang=’fr’). COMET uses the Unbabel/wmt22-comet-da model trained on human-annotated translations, which was shown to correlate strongly with human judgments in previous MT evaluation studies.
4.  Evaluated LLMs
@@
We evaluated 12 LLMs across different model families and scales. We evaluated GPT models from OpenAI – GPT-3.5-turbo, GPT-4, GPT-4-turbo, and GPT-4o – accessed through API requests. [^5] https://platform.openai.com We evaluated the following open-weight models: Mistral (7.25B), Phi-4 (14.7B), three LLaMA versions (3.1 with 8B, 3.2 with 3.2B, and 3.3 with 70B parameters), and three DeepSeek-R1 models (8B, 14B, and 32B). These were run locally using the Ollama framework, [^6] https://ollama.com on a Linux server with four NVIDIA RTX 2080 Ti GPUs (11 GB VRAM each), using parallel querying to maximize throughput. For each model, we used its default quantized versions provided by Ollama (Q4_0 or Q4_K_M). We do not report results with the Tower LLM (Alves et al., 2024), although it is fine-tuned specifically for translation, because it was not instructed for CoT prompting, and it obtained low scores in our pilot experiments on the contrastive task.

### v2

@@ Abstract
This paper assesses the ability of large language models (LLMs) to translate texts that include inter-sentential dependencies. We use the English-French DiscEvalMT benchmark (Bawden et al., 2018) with pairs of sentences containing translation challenges for pronominal anaphora and lexical cohesion. We evaluate 12 LLMs from the DeepSeek-R1, GPT, Llama, Mistral and Phi families on two tasks: (1) distinguish a correct translation from a wrong but plausible one; and (2) generate a correct translation. We compare prompts that encourage chain-of-thought reasoning with those that do not. The best models take advantage of reasoning and reach about 90% accuracy on the first task and COMET scores of about 92% on the second task, with GPT-4, GPT-4o and Phi standing out. Moreover, we observe a “wise get wiser” effect: the improvements through reasoning are larger for models that already perform well without reasoning.
Keywords: Context-aware MT, Translation with LLMs, Chain-of-thought reasoning for translation.
@@
Shabnam Ataee1, Hugo Huart1, Andrei Popescu-Belis1,2
@@
{shabnam.ataee, hugo.huart, andrei.popescu-belis}@heig-vd.ch
@@ 1.  Introduction
Large language models (LLMs) have shown impressive capacities for translation, among many other tasks, but their translations are still not perfect (Kocmi et al., 2024; Cui et al., 2025). To increase quality, we examine whether chain-of-thought (CoT) reasoning helps to improve referential and lexical cohesion across sentences. Indeed, when prompted to generate a step-by-step explanation of their answers, LLMs improve substantially on many tasks (Wei et al., 2022).
We hypothesize that CoT can improve the translation of elements that maintain coherence across sentences by making explicit translation decisions. To test our hypothesis, we use the DiscEvalMT English-French test suites for pronominal anaphora and lexical consistency designed by Bawden et al. (2018), in two settings. First, we ask an LLM to select the correct translation among the two contrastive alternatives present in the data. Second, we ask the LLM to translate each test sentence and we score the result against a reference translation. In both settings, we demonstrate the advantages of CoT prompting, though only for the largest LLMs, which improve their performance when encouraged to reason before translating.
The paper is organized as follows. We review related work in Section 2. We present the benchmark data and the evaluation metrics in Section 3. The tested LLMs are listed in Section 4: they include 4 GPT proprietary models from OpenAI and 8 open-weight models. The prompts and the results for the contrastive evaluation task (i.e., selecting the correct translation) are presented in Section 5, while those for the translation task are in Section 6. Our main contributions are the following ones:
@@
  • We show that the translation task is reliably scored, as BLEU, chrF, BERTscore, and COMET scores are correlated.
  • We demonstrate that CoT prompting improves coherence, though only for the best models. In other words, powerful models have a better capacity to leverage reasoning, which we call a “wise get wiser” effect.
  • We share the outputs of the tested systems at https://zenodo.org/records/17483104.
@@ 2.  Related Work
The translation capabilities of multilingual LLMs have attracted growing attention since 2023. Conversational LLMs are now often used for translation tasks, which represent around 4.5% of non-professional interactions with ChatGPT (Chatterji et al., 2025). Therefore, studies evaluating the translation capability of LLMs are of high importance, all the more that translation is included in common benchmarks for LLMs. LLMs have been tested since 2023 at the Workshops on Machine Translation (Kocmi et al., 2023, 2024), and in 2025, the majority of submissions to the general task were LLM-based (Kocmi et al., 2025). At the same time, the first studies of LLMs for MT included quantitative evaluations (Vilar et al., 2023; Zhang et al., 2023; Hendy et al., 2023). Some studies focused on evaluation only: for instance, Bawden and Yvon (2023) demonstrated the strong MT performance of BLOOM, particularly in few-shot settings and for high-resource language pairs. Fine-tuning of LLMs for document-level MT was found to be moderately beneficial (Wu et al., 2024).
@@
Word-sense disambiguation (WSD) for translation is one of the challenges on which LLMs have been evaluated. Test suites related to WSD include DiscEvalMT (Bawden et al., 2018) which we use in this paper and describe below, ContraWSD (Rios Gonzales et al., 2017) with about 7,000 sentences with reference translations, and a smaller test suite to score word translations (Rios et al., 2018). The DiBiMT benchmark with about 600 examples, for translation from English to five other languages, was proposed by Campolungo et al. (2022) but remains private and systems must be submitted to its owners. DiBiMT was used by Iyer et al. (2023) to score the capacity of LLMs for WSD.
@@
The capacity to translate potentially ambiguous pronouns can be tested using contrastive test suites such as DiscEvalMT (Bawden et al., 2018), ContraPRO (Müller et al., 2018; Lopes et al., 2020), or PROTEST (Guillou and Hardmeier, 2016), which was used at WMT 2018 (Guillou et al., 2018). A test suite for ellipses and lexical cohesion was designed by Voita et al. (2019). A review of scores reached on the DiscEvalMT data by various context-aware MT systems appears below in Section 5.2. An alternative approach is to align and compare pronouns in the MT output with those in a reference translation Miculicich Werlen and Popescu-Belis (2017), but scores may be biased by the choice of the antecedent.
@@
The correct translation of pronouns or ambiguous words is often related to the more general ability to leverage inter-sentential dependencies and generate a contextually-correct translation, reviewed in several studies (Popescu-Belis, 2019; Maruf et al., 2021; Jin et al., 2023; Castilho and Knowles, 2025). Document-level translation and its evaluation are important because MT outputs may appear competitive with human translations at the sentence level, but not at the document level (Läubli et al., 2018). An evaluation of GPT-3.5 and GPT-4 on document-level translation (Wang et al., 2023) found that they surpassed non-LLM systems in human ratings. Karpinska and Iyyer (2023) found that LLMs translate better entire paragraphs than individual sentences, in the case of literary translation. In an analysis of the translation of six types of contextual dependencies (with one sentence per type), Castilho et al. (2023) found that GPT 3.5 outperformed NMT systems on three high-resource languages, but not on a low-resource one (Gaelic). Beyond direct translation, interaction was a key element in the WSD process proposed by Pilault et al. (2023), using questions and answers as a chain-of-thought input to an LLM.
@@
Recent work has explored how reasoning can improve LLM-based translation, inspired by the seminal work of Wei et al. (2022), who introduced chain-of-thought (CoT) prompting and showed that guiding models through intermediate steps improves performance on complex tasks. Liu et al. (2025) argue that reasoning enhances translation by improving coherence, cultural alignment, and self-reflection. He et al. (2025) present R1-T1, which uses expert CoT templates and reinforcement learning to encourage inference-time reasoning for translation. Ye et al. (2025) found that reasoning models outperform standard LLMs on semantically complex and domain-specific MT, especially for longer texts.
From the variety of MT evaluation metrics, many have also been considered for extension to document-level evaluation (Dahan et al., 2024). Given that we only consider here sentence pairs, we will consider four widely used metrics. We will use BLEU and chrF (Papineni et al., 2002; Popović, 2015), which are based on surface-level overlap, although they offer limited insights into discourse-level quality. We will also use embedding-based metrics such as BERTScore (Zhang et al., 2020), and trained metrics such as COMET (Rei et al., 2020), as they model semantic similarity and are now frequently used.
@@
LLMs have also been used for MT evaluation: Kocmi and Federmann (2023) showed that GPT-based metrics reach state-of-the-art correlations with human judgments on WMT benchmarks. The benefits of LLM reasoning for this task are still under investigation: Larionov et al. (2025) found that reasoning-enabled LLMs such as DeepSeek-R1 and OpenAI’s o3-mini do not always outperform their non-reasoning counterparts in terms of alignment with human judgments of MT quality.
@@ 3.  Evaluation Data and Metrics
Anaphora Lexical choice
EN context It’s been a while since I last went to the river. And then the attack took place.
EN current It feels great to finally see it. A truly terrible attack.
FR context (given) Ça fait longtemps que je n’ai pas été à la rivière. Et puis l’attaque a eu lieu.
FR current #1 (correct) C’est chouette de la voir enfin. Une attaque vraiment affreuse.
FR current #2 (incorrect) C’est chouette de le voir enfin. Un assaut vraiment affreux.
Table 1: Contrastive test items for anaphora and lexical choice from DiscEvalMT (Bawden et al., 2018).
@@ 3.1.  Benchmark Data
This study uses the DiscEvalMT benchmark (Bawden, 2018; Bawden et al., 2018), designed to evaluate challenges in English-to-French translation that arise from two types of inter-sentential dependencies: pronominal anaphora and lexical cohesion. [^1] https://github.com/rbawden The test cases are manually constructed. Each test item consists of two English sentences (‘context’ and ‘current’), a French translation of the first one, and two alternative French translations of the second one: one is contextually appropriate, preserving coherence across the two sentences, while the other introduces a discourse-level error. Together, they form a contrastive pair among which a system should distinguish the correct translation.
An item for testing the translation of pronominal anaphora is shown in Table 1 (center). The word ‘river’ is translated as ‘rivière’ (feminine), so FR translation #1 has correct gender agreement between the object pronoun ‘la’ and the antecedent ‘rivière’, while the second one has wrong agreement. To control for chance, in another test item, ‘river’ is translated as ‘fleuve’ (masculine), making FR translation #2 correct. Two additional test items are made with approximate translations of the antecedent (e.g., ‘piscine’ (fem.) or ‘cinéma’ (masc.)). In these cases, the translation of the second sentence should still use a pronoun of the same gender as the antecedent, to maintain agreement.
For lexical choice, each item contains a word in the context sentence, which can be translated in several ways. The correct translation of the second sentence should use the same word as in the reference translation of the first sentence, to maintain consistency. In Table 1 (right), the word ‘attack’ is translated as ‘attaque’ in the context sentence, so the current sentence should reuse the same word, as in FR translation #1. In an additional test item, ‘attack’ is translated by ‘assaut’ in the context sentence, making FR translation #2 correct.
The anaphora test set includes 200 items (50 sets of four similar context sentences), each with two candidate translations of the second sentence. Similarly, the lexical choice test set includes 200 items with two candidates each. We selected the first half of the data for developing the prompts, and kept the second half for final testing.
@@ 3.2.  Metrics
In the first part of our study (Section 5), the task of the LLM is to distinguish the correct translation of the second sentence in each pair from the wrong one (contrastive task). In the second part (Section 6) the task of the LLM is to translate the second sentence into French.
@@
The contrastive task is scored by counting the number of times the correct translation was identified. The LLMs are prompted to output their response either as ‘Choice: (1)’ or as ‘Choice: (2)’. This can be preceded by reasoning in case of CoT prompting. Our evaluation script is tolerant to minor formatting differences in the LLM’s output: responses are accepted as correct as long as they clearly indicate the right option (1 or 2); if they cannot be parsed, they are counted as incorrect. The mean accuracy score is the average over all the answers.
@@
To avoid systematic bias in favor of either candidate #1 or #2, we present each test item twice, once with the correct option being first, and once second. Accordingly, we measure consistency, i.e. the sensitivity of the model to the position of the correct translation among the two options. Ideally, the response should not depend on the position. We define inconsistency as the absolute difference between the accuracy score when the correct option is presented first ($\text{ACC}_{\text{correct=1}}$) and the accuracy score when it is presented second ($\text{ACC}_{\text{correct=2}}$), normalized by the sum of these accuracies. Inconsistency varies from 0, if accuracy is not influenced at all by the position of the correct answer, to 1, e.g. if an LLM always answers ‘1’.
$\text{Inc.}=\frac{|\text{ACC}_{\text{correct=1}}-\text{ACC}_{\text{correct=2}}|}{\text{ACC}_{\text{correct=1}}+\text{ACC}_{\text{correct=2}}}$
@@
For the translation task, we provide the LLM with the two EN source sentences, as well as the FR translation of the context sentence, and ask it to translate the second EN sentence (possibly outputting reasoning). We parse the output to isolate the translation, and score it by comparing it to the correct translation using four metrics: BLEU, chrF, BERTscore and COMET. The first two capture surface-level overlap, while BERTScore and COMET assess semantic similarity using embeddings. All metrics are computed using their official Python implementations available via the sacrebleu [^2] https://github.com/mjpost/sacrebleu, bert-score [^3] https://github.com/Tiiiger/bert_score, and unbabel-comet [^4] https://github.com/Unbabel/COMET packages. BLEU and chrF are calculated with the default configurations of sacrebleu. BERTScore is computed using the default multilingual model (bert-base-multilingual-cased) and the language parameter set to French (lang=’fr’). COMET uses the Unbabel/wmt22-comet-da model trained on human-annotated translations, which was shown to correlate strongly with human judgments. Given the high density of pronouns, we believe that scoring the entire sentence with frequently-used global metrics is a better solution than comparing pronouns with the APT metric (Miculicich Werlen and Popescu-Belis, 2017).
4.  Evaluated LLMs and Baseline
@@
We evaluated 12 LLMs across different model families and scales. We evaluated GPT models from OpenAI – GPT-3.5-turbo, GPT-4, GPT-4-turbo, and GPT-4o – accessed through API requests. [^5] https://platform.openai.com We evaluated the following open-weight models: Mistral (7.25B), Phi-4 (14.7B), three versions of Llama (3.1 with 8B, 3.2 with 3.2B, and 3.3 with 70B parameters), and three versions of DeepSeek-R1 (8B, 14B, and 32B); the latter are trained to reason by default, as discussed in Section 5 below. We do not report results with the Tower LLM (Alves et al., 2024), although it is a model that was fine-tuned specifically for translation. However, Tower was not instructed for CoT prompting, and we obtained low scores with it in preliminary experiments.
The LLMs were run locally using the Ollama framework, [^6] https://ollama.com which provides its own versions of the above models. For each model, we used the default quantized versions provided by Ollama (Q4_0 or Q4_K_M). Experiments were done on a Linux server with four NVIDIA RTX 2080 Ti GPUs with 11 GB VRAM each, using parallel querying to maximize throughput, with up to 16 simultaneous query threads. This limit was determined empirically to ensure stable and efficient throughput.
We include for comparison an encoder-decoder neural MT system, namely the distilled NLLB-200 multilingual model with 600M parameters (NLLB-Team et al., 2022). The model was queried using the Transformers library from Hugging Face. [^7] https://huggingface.co/facebook/nllb-200-distilled-600M. For both tasks (contrastive and translation), we used prefix decoding: the French translation of the first sentence was added as a fixed prefix to the decoder. Then, for the contrastive task, we generate the translation of the second sentence, and compare it to each of the two options in the test pair using BERTScore. We assume that the most similar one is the model’s answer. For the translation task, the MT model also uses prefix decoding to generate the translation of the second sentence, which is then evaluated.